QUESTION Which of Which of the tollowing wouild rsut in the highest annual income blo he riger arr al roome lassurng a 40-hour workweek and 52 weeks tr positional? yho of rig woud-sut
Answer #1

1. Option A has highest income.

A. Income = $18*40*52 = $ 37,440 per year

B. Income = $ 25,000 per year

C. Income = $1,900*12 = $ 22,800 per year

D. Income = $ 650 *52 = $ 33,800 per year

E. Income = $ 1350*(52/2) = $ 35,100

2. Highest income = $ 4200 per month

  • Annual income = $1750*26 = $ 45,500
  • Annual income = $900*52 = $ 46,800
  • Annual income = $ 22*40*52 = $ 45,760
  • Annual income =$4200*12 = $ 50,400
  • Annual income = $ 50,000

3. Highest income = $ 12 per hour

  • Annual income =$450*52 =$ 23,400
  • Annual income = $ 750*26 = $ 19,500
  • Annual income = $ 1200*12 = $ 14,400
  • Annual income = $ 20,000
  • Annual income =$12*40*52 = $ 24,960

4. Highest income = $ 3000 per month

  • Annual income = $ 625*52 = $ 32,500
  • Annual income = $ 1200*26 = $ 31,200
  • Annual income = $ 30,000
  • Annual income = $12*40*52 = $ 24,960
  • Annual income = $3000*12 = $ 36,000

5. Highest income =$ 3,200 per month

  • Annual income = $ 3,200*12 = $ 38,400
  • Annual income = $ 17*40*52 =$ 35,360
  • Annual income = $ 40,000
  • Annual income = $ 1450*26 = $ 37,700
  • Annual income = $ 710*52 = $ 36,920

Please contact if having any query thank you.